NvidiaIntely (nvidiaintely):

A balloon contains a fixed amount of helium. If the initial conditions are 1.12 liters, 3.00 atmospheres, and 19 degrees Celsius, what will the final volume be when the balloon is exposed to 1.00 atmosphere pressure and 22 degrees Celsius? 0.370 L 0.377 L 2.90 L 3.39 L

NvidiaIntely (nvidiaintely):

@Google I believe its 3.39 L.

OpenStudy (google):

hmm

OpenStudy (google):

hold on gimme a minute

NvidiaIntely (nvidiaintely):

Okay (;

OpenStudy (google):

you know how to do the math on this question?

NvidiaIntely (nvidiaintely):

V1 = 1.12 L P1 = 3 atm T1 = 19 + 273 = 292 K V2 = ? P2 =1 atm T2 = 22 + 273 = 295 k P1V1/T1 = P2V2/T2 V2 = P1V1T2/P2T1 = (3)(1.12)(295)/(1)292) = 3.39 Liters So my answer = (D) 3.39 L

OpenStudy (google):

\[3.00x1.12/ (19+273) = 1.00 x V2 / (22+273) \]

OpenStudy (google):

your answer is right but the math is kinda off

OpenStudy (google):

\[v2=3.93L \]

OpenStudy (google):

is correct

NvidiaIntely (nvidiaintely):

Okay, got ya.